# BPINN-EM-Post: Bayesian Physics-Informed Neural Network based Stochastic Electromigration Damage Analysis in the Post-void Phase

### v2

@@ I Introduction
Traditionally, the MTTF or reliability of an interconnect wire is linked to the wire’s current density through well-established EM models by Black and Blech [2, 3]. However, these single-wire segment EM models have been increasingly criticized for being overly conservative, as they treat each wire segment in isolation. Recent research demonstrates that stress evolution among wire segments within an EM tree, is highly correlated and should be analyzed collectively [4, 5]. Numerous physics-based analytical and numerical solutions have been proposed to address the above-mentioned limitations [6, 7, 8, 9, 10, 11, 12, 13, 14]. These methods primarily focus on solving Korhonen’s partial differential equation (PDE) to model EM-induced stress evolution [15]. Nevertheless, traditional numerical techniques for solving Korhonen’s equation and related PDEs remain computationally costly. Furthermore, stochastic physics-based EM analyses using Monte Carlo simulations are prohibitively expensive for numerical solvers such as COMSOL [16] and EMSpice [13], which demand multiple samplings to estimate variation distributions.
@@
In recent years, a machine learning-based method known as physics-informed neural networks (PINN) has been developed to facilitate the learning and integration of physical laws represented by nonlinear PDEs in complex physical, biological, and engineering systems [17, 18]. PINN approach has been utilized to solve Korhonen’s equation, showing promising results for EM stress modeling [19, 20, 21]. In this framework, physical laws, boundary conditions (BCs), and initial conditions (ICs) of the PDEs are explicitly enforced through loss functions within the neural networks. This approach has demonstrated effectiveness on small-scale PDE problems. Recently, hierarchical PINN schemes have been introduced to accelerate EM analysis [20, 22]. These approaches divide the stress evolution problem in multi-segment interconnects into two stages, thereby significantly reducing the number of variables, as well as enhancing computational efficiency and scalability to large-scale interconnect structures. However, these PINN-based methods have yet to tackle the variability in EM effects.
@@
Bayesian Neural Networks (BNNs) have emerged as an important class of neural networks due to their robust capabilities in uncertainty quantification, handling limited datasets, and preventing overfitting [23]. Lamichhane et al. introduced a hierarchical approach that integrates a pre-trained Bayesian network with a PINN framework to estimate EM stress variations, with a particular focus on the nucleation phase [24]. However, extending BNNs to analyze the post-voiding phase poses significant challenges, as it requires incorporating the initial stress distribution, a task that is highly complex and often infeasible for even simple neural networks [22].
@@ II-A EM Stress Evolution in Post-voiding Phase
In confined metal wires subjected to high current densities, EM occurs as atoms are migrated from the cathode to the anode due to interactions between electrons and metal atoms [2]. Over time, this atom migration may potentially cause void and hillock formulations that compromise interconnect functionality. Various methods have been developed to evaluate the EM-induced reliability of metal interconnects. Among these, Blech’s limit [3] and Black’s MTTF [2] are traditional approaches for reliability assessment. Despite their utility, these methods are becoming less popular as Black’s MTTF is limited to individual line segments and neither method accounts for the evolution of hydrostatic stress. Korhonen’s equation [15] addresses these limitations by describing EM in confined metal wires through a set of PDEs, and extends to model hydrostatic stress in multi-segment interconnects [25, 12]. The equation for the nucleation phase in multi-segment interconnects is given by Eq. (1), where $\sigma_{ij}(x,t)$ represents the stress in the interconnect segment $ij$ connecting nodes $i$ and $j$. In Eq. (1), $G_{ij}$ denotes the EM driving force in segment $ij$, calculated as $G_{ij}=\frac{e\rho J_{ij}Z^{*}}{\Omega}$, where $J_{ij}$ is the current density in segment $ij$.
@@
For analyzing multi-segment interconnects during the post-voiding phase, Eq. (LABEL:eq:korhonen_equation_single_wire_postvoid) is applied to the segment that contains the void. Segments without voids are addressed using Eq. (1). To accurately determine the stress profile across the entire interconnect tree, BCs that ensure stress continuity and atomic flux conservation must be implemented [27].
@@ II-B Estimation of Variations in EM Stress Evolution
here $\bar{\sigma}$ represents the EM stress outcomes calculated using the parameter set $\mathbf{x}$ with the addition of noise $v$ from variability. $\mathcal{F}(.)$ could be any solver employed to solve combination of Eq. (1) and Eq. (LABEL:eq:korhonen_equation_single_wire_postvoid), such as the FEM-based COMSOL method [16], or the FDM-based EMSpice tool [13]. Variation noise is typically assumed to be independent and normally distributed with a mean of zero and a specified variance $Var_{\sigma}$ [28]. A common technique for estimating the mean and variance in EM stress calculations, as described by combinations of Eq. (1) and Eq. (LABEL:eq:korhonen_equation_single_wire_postvoid), involves direct Monte Carlo simulation. Nonetheless, Monte Carlo simulations necessitate repeated runs of combination of Eq. (1) and Eq. (LABEL:eq:korhonen_equation_single_wire_postvoid) for multi-segment interconnects, which can become exceedingly expensive, particularly as the number of segments in the interconnect structure grows.
@@ III-A Estimation of Prior for BPINN
The overall framework of our proposed BPINN-EM-Post is shown in Fig. 1. A fully connected network serves as a surrogate model for our Bayesian physics-informed neural network (BPINN) framework, based on prior work [20, 24] and the demonstrated capability of Multi-layer Perceptron (MLP) architecture for this prediction task.
@@
The unknown parameters $\omega$ comprise all weight matrices $\mathbf{w}$ and bias vectors $\mathbf{b}$. For the BNN, $\omega$ is assigned a prior such that its components are independent Gaussian distributions with mean 0 and variances $\mathbf{Var}_{w,l}$ and $\mathbf{Var}_{b,l}$ for all layers [33]. As the hidden layer width approaches infinity, $\tilde{\mathbf{F}}(\mathbf{x})$ converges to a Gaussian process [33].
@@ III-B Generating Observations with Analytical Solutions
To generate stress solutions $\tilde{\sigma}(x,t;\omega)$ using $\tilde{\mathbf{F}}(\mathbf{x};\omega)$, we adapt the analytical solutions proposed in [22]. Before using the output from the Bayesian network, $\tilde{\mathbf{F}}(\mathbf{x};\omega)$, in the analytical solution, we perform linear transformations on it.
@@
Using flux information at each inter-segment junction, we calculate flux values $F^{-}$ and $F^{+}$ at the endpoints of each segment within the multisegment interconnect structure. The analytical solutions in PostPINN-EM [22] use these endpoint flux values to compute stress on single wire segments. Here, $F^{-}$ and $F^{+}$ represent flux at the left/down and right/upper positions, respectively.
@@
For the segments with void, we assume that one end of the segment could either be an inter-segment junction or a terminal boundary, while the other end corresponds to a void node. Fig. 2(b) illustrates a single segment with a void at the $x=L$ end. At $x=0$, which may represent either an inter-segment junction or a terminal boundary, a modifiable BC from the voidless scenario is applied. We modify the BC at $x=L$ to $\tilde{\sigma}(L,t)=0$. When using a very small $\delta$ in Eq. (LABEL:eq:korhonen_equation_single_wire_postvoid), the BC at the site of the void mirrors the approach in [34]. The BCs for the segment with a void at the subsequent node of a wire segment, i.e., at $x=L$, are expressed as Eq. (9),
@@
Similarly, when the void is positioned at the preceding (left/bottom) end, i.e., at $x=0$, the BC is defined as $\tilde{\sigma}(x,t)=0$ at $x=0$ and $\frac{\partial\tilde{\sigma}(x,t)}{\partial x}=\phi^{+}(t)$ at $x=L$. Using the same Laplace transform method, the analytical solution for this scenario can also be derived. In Eq. (7) and Eq. (10) the values $\phi^{-}(0)$ and $\phi^{+}(0)$ can be calculated using the current density information at each inter-segment junctions [22, 21]. For each single segment, we use the flux information from Bayesian network as $\frac{d\phi^{-}(t)}{dt}=F^{-}(\mathbf{x};\omega)$ and $\frac{d\phi^{+}(t)}{dt}=F^{+}(\mathbf{x};\omega)$.
@@ III-C Evaluation of Posterior for BPINN
Eq. (19) is discretized through the leapfrog method, and the Metropolis-Hastings step is employed to minimize the discretization error. More implementation details on the HMC algorithm are explained in [33].
@@ IV-B Data Preparation and Scaling
To determine the stress distribution in a multi-segment interconnect structure, we first gather the physical and material attributes of the interconnect segments. Each segment $i$ is characterized by its physical properties, including length $L_{i}$ and width $w_{i}$, while assuming uniform thermal conductivity $\kappa$ across all segments. Spatial points $x$ are sampled within $[0,L]$, and temporal points are chosen from $t=0$ to $1\times 10^{9}$ s at intervals of $1\times 10^{7}$ s. In this analysis, we incorporate the nucleation-phase solutions for these interconnects that are easily pre-established by existing methods [13, 16, 24] consistent to this work, providing initial stress distribution $h(x)$ and void locations.
@@
TABLE I: Performance and accuracy comparisons with existing methods [16][13]
# of wires COMSOL [16] EMSpice [13] The Proposed BPINN-EM-Post
@@
To assess the accuracy and performance of our proposed BPINN-EM-Post, we compare results against Monte Carlo simulations using FEM-based COMSOL [16] and FDM-based EMSpice tool [13]. For each Monte Carlo iteration, branch currents are randomly sampled from a normal distribution with a mean range of $5\times 10^{-8}\,\mathrm{A/m^{2}}$ to $5\times 10^{9}\,\mathrm{A/m^{2}}$ to introduce variability to input currents.
@@ IV-D Overall Accuracy and Performance Analysis
To assess the efficacy and precision of our method across a variety of general multi-segment interconnect structures, we analyzed a total of 1000 randomly generated configurations, each containing between 50 and 250 segments, with each inter-segment junction connected to at most four segments. These structures were examined to gauge variational EM stress through the use of COMSOL [16], EMSpice [13], and our proposed BPINN-EM-Post, utilizing 30 samples to measure variations in stochastic EM stress. We should notice that other existing works [20, 22] are not applicable for stochastic analysis, while [24] cannot be employed in post-voiding phase, thus these methods are not available for numerical comparison. The findings, detailing accuracy and performance metrics for multi-segment interconnect structures across different segment counts, are showcased in Tab. I. Our method exhibits superior acceleration than prior arts, with error rate of only 1.10% at most, compared with COMSOL and less than 0.2% compared with EMSpice. In addition, Training and Inference time shows nearly linear growth with respect to the number of wires, indicating its scalability to larger interconnect structures.
@@
Fig. 4 demonstrates the variations in EM stress, expressed through both mean and standard deviation, for a ten-segment interconnect structure for simplicity. The physical and electrical characteristics of which are shown in Fig. 4(a). The variations in EM stress at the selected junctions between segments of this interconnect structure are detailed in Fig. 4(b), Fig. 4(c), Fig. 4(d), and Fig. 4(e). In these comparisons, our proposed BPINN-EM-Post is shown to precisely predict the variations in EM stress, aligning closely with the results obtained from COMSOL [16] and EMSpice [13].

### v3

@@ I Introduction
Traditionally, the MTTF or reliability of an interconnect wire is linked to the wire’s current density through well-established EM models by Black and Blech [Black:1969fc, Blech:1976ko]. However, these single-wire segment EM models have been increasingly criticized for being overly conservative, as they treat each wire segment in isolation. Recent research demonstrates that stress evolution among wire segments within an EM tree, is highly correlated and should be analyzed collectively [5, 15]. Numerous physics-based analytical and numerical solutions have been proposed to address the above-mentioned limitations [DeORIO:Micro'2010, 14, ChenTan:TDMR'17, 3, CookSun:TVSI'18, Abbasinasab:DAC'2018, ChenTan:TVLSI'19, SunYu:TDMR'20, 1, Lamichhane:SMACD'25]. These methods primarily focus on solving Korhonen’s partial differential equation (PDE) to model EM-induced stress evolution [Korhonen:1993bb]. Nevertheless, traditional numerical techniques for solving Korhonen’s equation and related PDEs remain computationally costly. Furthermore, stochastic physics-based EM analyses using Monte Carlo simulations are prohibitively expensive for numerical solvers such as COMSOL [comsol:heat'2014] and EMSpice [SunYu:TDMR'20], which demand multiple samplings to estimate variation distributions.
@@
In recent years, machine learning–based methods have demonstrated strong effectiveness and efficiency in addressing complex EDA problems [SadiqbatchaZhao:ASPDAC'20, Jin:ICCAD'20, Lu:ICCAD'23, Lu:ISLPED'25]. Among them, physics-informed neural networks (PINN) has been developed to facilitate the learning and integration of physical laws represented by nonlinear PDEs in complex physical, biological, and engineering systems [13, 12]. PINN approach has been utilized to solve Korhonen’s equation, showing promising results for EM stress modeling [6, Jin:ICCAD'22, 7]. In this framework, physical laws, boundary conditions (BCs), and initial conditions (ICs) of the PDEs are explicitly enforced through loss functions within the neural networks. This approach has demonstrated effectiveness on small-scale PDE problems. Recently, hierarchical PINN schemes have been introduced to accelerate EM analysis [Jin:ICCAD'22, Lamichhane:ICCAD'23]. These approaches divide the stress evolution problem in multi-segment interconnects into two stages, thereby significantly reducing the number of variables, as well as enhancing computational efficiency and scalability to large-scale interconnect structures. However, these PINN-based methods have yet to tackle the variability in EM effects.
@@
Bayesian Neural Networks (BNNs) have emerged as an important class of neural networks due to their robust capabilities in uncertainty quantification, handling limited datasets, and preventing overfitting [4]. Lamichhane et al. introduced a hierarchical approach that integrates a pre-trained Bayesian network with a PINN framework to estimate EM stress variations, with a particular focus on the nucleation phase [Lamichhane:ICCAD'24]. However, extending BNNs to analyze the post-voiding phase poses significant challenges, as it requires incorporating the initial stress distribution, a task that is highly complex and often infeasible for even simple neural networks [Lamichhane:ICCAD'23].
@@ II-A EM Stress Evolution in Post-voiding Phase
In confined metal wires subjected to high current densities, EM occurs as atoms are migrated from the cathode to the anode due to interactions between electrons and metal atoms [Black:1969fc]. Over time, this atom migration may potentially cause void and hillock formulations that compromise interconnect functionality. Various methods have been developed to evaluate the EM-induced reliability of metal interconnects. Among these, Blech’s limit [Blech:1976ko] and Black’s MTTF [Black:1969fc] are traditional approaches for reliability assessment. Despite their utility, these methods are becoming less popular as Black’s MTTF is limited to individual line segments and neither method accounts for the evolution of hydrostatic stress. Korhonen’s equation [Korhonen:1993bb] addresses these limitations by describing EM in confined metal wires through a set of PDEs, and extends to model hydrostatic stress in multi-segment interconnects [WangYan:ICCAD'17, ChenTan:TVLSI'19]. The equation for the nucleation phase in multi-segment interconnects is given by Eq. (1), where $\sigma_{ij}(x,t)$ represents the stress in the interconnect segment $ij$ connecting nodes $i$ and $j$. In Eq. (1), $G_{ij}$ denotes the EM driving force in segment $ij$, calculated as $G_{ij}=\frac{e\rho J_{ij}Z^{*}}{\Omega}$, where $J_{ij}$ is the current density in segment $ij$.
@@
For analyzing multi-segment interconnects during the post-voiding phase, Eq. (LABEL:eq:korhonen_equation_single_wire_postvoid) is applied to the segment that contains the void. Segments without voids are addressed using Eq. (1). To accurately determine the stress profile across the entire interconnect tree, BCs that ensure stress continuity and atomic flux conservation must be implemented [Huang:TCAD'15].
@@ II-B Estimation of Variations in EM Stress Evolution
here $\bar{\sigma}$ represents the EM stress outcomes calculated using the parameter set $\mathbf{x}$ with the addition of noise $v$ from variability. $\mathcal{F}(.)$ could be any solver employed to solve combination of Eq. (1) and Eq. (LABEL:eq:korhonen_equation_single_wire_postvoid), such as the FEM-based COMSOL method [comsol:heat'2014], or the FDM-based EMSpice tool [SunYu:TDMR'20]. Variation noise is typically assumed to be independent and normally distributed with a mean of zero and a specified variance $Var_{\sigma}$ [8]. A common technique for estimating the mean and variance in EM stress calculations, as described by combinations of Eq. (1) and Eq. (LABEL:eq:korhonen_equation_single_wire_postvoid), involves direct Monte Carlo simulation. Nonetheless, Monte Carlo simulations necessitate repeated runs of combination of Eq. (1) and Eq. (LABEL:eq:korhonen_equation_single_wire_postvoid) for multi-segment interconnects, which can become exceedingly expensive, particularly as the number of segments in the interconnect structure grows.
@@ III-A Estimation of Prior for BPINN
The overall framework of our proposed BPINN-EM-Post is shown in Fig. 1. A fully connected network serves as a surrogate model for our Bayesian physics-informed neural network (BPINN) framework, based on prior work [Jin:ICCAD'22, Lamichhane:ICCAD'24] and the demonstrated capability of Multi-layer Perceptron (MLP) architecture for this prediction task.
@@
The unknown parameters $\omega$ comprise all weight matrices $\mathbf{w}$ and bias vectors $\mathbf{b}$. For the BNN, $\omega$ is assigned a prior such that its components are independent Gaussian distributions with mean 0 and variances $\mathbf{Var}_{w,l}$ and $\mathbf{Var}_{b,l}$ for all layers [Yang:JCP'19]. As the hidden layer width approaches infinity, $\tilde{\mathbf{F}}(\mathbf{x})$ converges to a Gaussian process [Yang:JCP'19].
@@ III-B Generating Observations with Analytical Solutions
To generate stress solutions $\tilde{\sigma}(x,t;\omega)$ using $\tilde{\mathbf{F}}(\mathbf{x};\omega)$, we adapt the analytical solutions proposed in [Lamichhane:ICCAD'23]. Before using the output from the Bayesian network, $\tilde{\mathbf{F}}(\mathbf{x};\omega)$, in the analytical solution, we perform linear transformations on it.
@@
Using flux information at each inter-segment junction, we calculate flux values $F^{-}$ and $F^{+}$ at the endpoints of each segment within the multisegment interconnect structure. The analytical solutions in PostPINN-EM [Lamichhane:ICCAD'23] use these endpoint flux values to compute stress on single wire segments. Here, $F^{-}$ and $F^{+}$ represent flux at the left/down and right/upper positions, respectively.
@@
For the segments with void, we assume that one end of the segment could either be an inter-segment junction or a terminal boundary, while the other end corresponds to a void node. Fig. 2(b) illustrates a single segment with a void at the $x=L$ end. At $x=0$, which may represent either an inter-segment junction or a terminal boundary, a modifiable BC from the voidless scenario is applied. We modify the BC at $x=L$ to $\tilde{\sigma}(L,t)=0$. When using a very small $\delta$ in Eq. (LABEL:eq:korhonen_equation_single_wire_postvoid), the BC at the site of the void mirrors the approach in [Bondurant_2005]. The BCs for the segment with a void at the subsequent node of a wire segment, i.e., at $x=L$, are expressed as Eq. (9),
@@
Similarly, when the void is positioned at the preceding (left/bottom) end, i.e., at $x=0$, the BC is defined as $\tilde{\sigma}(x,t)=0$ at $x=0$ and $\frac{\partial\tilde{\sigma}(x,t)}{\partial x}=\phi^{+}(t)$ at $x=L$. Using the same Laplace transform method, the analytical solution for this scenario can also be derived. In Eq. (7) and Eq. (10) the values $\phi^{-}(0)$ and $\phi^{+}(0)$ can be calculated using the current density information at each inter-segment junctions [Lamichhane:ICCAD'23, 7]. For each single segment, we use the flux information from Bayesian network as $\frac{d\phi^{-}(t)}{dt}=F^{-}(\mathbf{x};\omega)$ and $\frac{d\phi^{+}(t)}{dt}=F^{+}(\mathbf{x};\omega)$.
@@ III-C Evaluation of Posterior for BPINN
Eq. (19) is discretized through the leapfrog method, and the Metropolis-Hastings step is employed to minimize the discretization error. More implementation details on the HMC algorithm are explained in [Yang:JCP'19].
@@ IV-B Data Preparation and Scaling
To determine the stress distribution in a multi-segment interconnect structure, we first gather the physical and material attributes of the interconnect segments. Each segment $i$ is characterized by its physical properties, including length $L_{i}$ and width $w_{i}$, while assuming uniform thermal conductivity $\kappa$ across all segments. Spatial points $x$ are sampled within $[0,L]$, and temporal points are chosen from $t=0$ to $1\times 10^{9}$ s at intervals of $1\times 10^{7}$ s. In this analysis, we incorporate the nucleation-phase solutions for these interconnects that are easily pre-established by existing methods [SunYu:TDMR'20, comsol:heat'2014, Lamichhane:ICCAD'24] consistent to this work, providing initial stress distribution $h(x)$ and void locations.
@@
TABLE I: Performance and accuracy comparisons with existing methods [comsol:heat'2014][SunYu:TDMR'20]
# of wires COMSOL [comsol:heat'2014] EMSpice [SunYu:TDMR'20] The Proposed BPINN-EM-Post
@@
To assess the accuracy and performance of our proposed BPINN-EM-Post, we compare results against Monte Carlo simulations using FEM-based COMSOL [comsol:heat'2014] and FDM-based EMSpice tool [SunYu:TDMR'20]. For each Monte Carlo iteration, branch currents are randomly sampled from a normal distribution with a mean range of $5\times 10^{-8}\,\mathrm{A/m^{2}}$ to $5\times 10^{9}\,\mathrm{A/m^{2}}$ to introduce variability to input currents.
@@ IV-D Overall Accuracy and Performance Analysis
To assess the efficacy and precision of our method across a variety of general multi-segment interconnect structures, we analyzed a total of 1000 randomly generated configurations, each containing between 50 and 250 segments, with each inter-segment junction connected to at most four segments. These structures were examined to gauge variational EM stress through the use of COMSOL [comsol:heat'2014], EMSpice [SunYu:TDMR'20], and our proposed BPINN-EM-Post, utilizing 30 samples to measure variations in stochastic EM stress. We should notice that other existing works [Jin:ICCAD'22, Lamichhane:ICCAD'23] are not applicable for stochastic analysis, while [Lamichhane:ICCAD'24] cannot be employed in post-voiding phase, thus these methods are not available for numerical comparison. The findings, detailing accuracy and performance metrics for multi-segment interconnect structures across different segment counts, are showcased in Tab. I. Our method exhibits superior acceleration than prior arts, with error rate of only 1.10% at most, compared with COMSOL and less than 0.2% compared with EMSpice. In addition, Training and Inference time shows nearly linear growth with respect to the number of wires, indicating its scalability to larger interconnect structures.
@@
Fig. 4 demonstrates the variations in EM stress, expressed through both mean and standard deviation, for a ten-segment interconnect structure for simplicity. The physical and electrical characteristics of which are shown in Fig. 4(a). The variations in EM stress at the selected junctions between segments of this interconnect structure are detailed in Fig. 4(b), Fig. 4(c), Fig. 4(d), and Fig. 4(e). In these comparisons, our proposed BPINN-EM-Post is shown to precisely predict the variations in EM stress, aligning closely with the results obtained from COMSOL [comsol:heat'2014] and EMSpice [SunYu:TDMR'20].